OpenStudy (anonymous):

what is the standard form of the equation that has a slope of (2/3) and passes through the point (1,0)

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

Slope of line is y = mx + b ; m = slope So, y = 2/3x + b You have to find b. You know the line passes through (1,0) so x = 1 and y = 0 So 0 = 2/3(1) + b b = -2/3 So your equation is y = 2/3 x - 2/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its either 3x-2y=2

OpenStudy (anonymous):

I don't understand your last comment.